find the area outside circle r = 1/2 and inside circle r = cosθ



1
Expert's answer
2021-10-27T15:42:29-0400

The two curves intersect where "1/2=\cos \theta." So "\theta=-\pi/3" or "\pi/3."

The area we want is then


"\dfrac{1}{2}\displaystyle\int_{-\pi/3}^{\pi/3}((\cos \theta)^2-(1/2)^2)d\theta"


"=\dfrac{1}{8}\displaystyle\int_{-\pi/3}^{\pi/3}(2(1+\cos(2\theta))-1)d\theta"

"=\dfrac{1}{8}\displaystyle\int_{-\pi/3}^{\pi/3}(1+2\cos(2\theta))d\theta"

"=\dfrac{1}{8}[\theta+\sin(2\theta)]\begin{matrix}
 \pi/3 \\
 -\pi/3
\end{matrix}"


"=\dfrac{1}{8}(\dfrac{\pi}{3}+\sin(\dfrac{2\pi}{3})-(-\dfrac{\pi}{3})-\sin(\dfrac{-2\pi}{3}))"

"=(\dfrac{\pi}{12}+\dfrac{\sqrt{3}}{8})({units}^2)"

The area outside circle "r = 1\/2" and inside circle "r = \cosθ" is "(\dfrac{\pi}{12}+\dfrac{\sqrt{3}}{8})" square units.